OpenStudy (anonymous):

A sample of a compound of Cl and O reacts with an excess of H2 to give 0.233 g of HCl and 0.403 g of H2O. Determine the empirical formula of the compound.

OpenStudy (dtan5457):

You still here?

OpenStudy (dtan5457):

All you need to do here is calculate the moles. For CL, it is given as 0.233, but the formula mass for HCL is 36.46. 0.233/36.46=0.00639 moles For O, it is given as 0.403, but the formula mass is 18.02. 0.403/18.02=0.0024 moles now just divide both by the smallest number 0.00639/0.00639=1 0.0024/0.00639=3.5 you can't have decimals, so multiply both by 2. Your empirical formula is CL2O7

OpenStudy (anonymous):

thanks